OpenStudy (anonymous):

-6x - 3y= 3 -3x + 5y = -5 A. (1,-1) b. (0,0) c. (0,-1) D. (1,0)

OpenStudy (anonymous):

you want to check the answers given, or just go ahead and solve?

OpenStudy (anonymous):

just go ahead and solve its my daughters paper i just am checking the right answer

OpenStudy (anonymous):

should be pretty clear that \((0,0)\) doesn't work, since that would give you \(0\) and not \(3\)

OpenStudy (anonymous):

yea but which one is the corect answer

OpenStudy (anonymous):

or solve \[-6x - 3y= 3\\ -3x + 5y = -5 \] by multiplying the second equation by \(-2\) to get the system \[-6x-3y=3\\6x-10y=10\]

OpenStudy (anonymous):

add the two equations (in a column) and get \[-13y=13\] which tells you \(y=-1\)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

put \(-1\) in for \(y\) in either equation, and you will find that \(x=0\)

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

for example \[-6x - 3y= 3\\-6x-3\times -1=3\\-6x+3=3\\-6x=0\\x=0\]

OpenStudy (anonymous):

yw